Homework: Chapter 9 -Continuing Problem Score: 0 of 10 pts CP943 (similar to) Estimated Acquisition Life SL any All Please answer the monthly depreciation expense and explain it. Thank you
Answer #1

Answers

Monthly depreciation expense

Computer

$250

Furniture

$35

Calculations

Double declining Method—Computer

A

Cost

$   9,000.00

B

Residual Value

$       900.00

C=A - B

Depreciable base

$   8,100.00

D

Life [in years]

6

E=C/D

Annual SLM depreciation

$   1,350.00

F=E/C

SLM Rate

16.67%

G=F x 2

DDB Rate

33.33%

H=(G x A)/12

Depreciation for march

$ 250.00

Straight line Method – Furniture

A

Cost

$   5,800.00

B

Residual Value

$   1,600.00

C=A - B

Depreciable base

$   4,200.00

D

Life [in years left ]

                  10

E=C/D

Annual SLM depreciation

$       420.00

F=E/12

Monthly depreciation

$         35.00